OpenStudy (anonymous):

The unit vector oppositely directed to −4i+j is??? i need help please.

OpenStudy (anonymous):

vector is <-4,1> for opposite direction multiply with -1 <4,-1> magnitude is \[\Large \sqrt{4^2+1^2}=\sqrt{17}\] divide each component with magnitude to get unit vector.

OpenStudy (anonymous):

i tried <4,-1>. its wrong

OpenStudy (anonymous):

the required answer will be when you divide by magnitude \[\Large <\frac{4}{\sqrt{17}},\frac{-1}{\sqrt{17}}>\]

OpenStudy (anonymous):

could u also solve The unit vector that has the same direction as the vector from the point A=(−2,1) to the point B=(−4,2) is??